# Towards Evaluating and Building Versatile Large Language Models for Medicine

### v1

@@ 2.3 Quantitative Results on Various Tasks
InternLM 2 achieves the highest accuracy on PubMedQA Answer Verification and HealthFact Verification, with scores of 98 and 92, respectively. In the EBMS benchmark, GPT-4 demonstrates the strongest performance, with BLEU/ROUGE scores of 16.36/16.33. Claude-3.5 follows closely with scores of 14.22/15.82, though it performs poorly on PubMedQA Answer Verification. Llama 3 achieves accuracy of 94 and 64 on PubMedQA and HealthFact Verification, respectively, and a BLEU/ROUGE score of 12.96/14.37. Mistral, while achieving a higher BLEU/ROUGE score of 15.19/15.56, has relatively lower accuracy of 54 and 68. MEDITRON ranks lowest across all metrics, with accuracy of 26 and 40 and BLEU/ROUGE scores of 14.08/25.69.
MMedIns-Llama 3 continues surpassing existing models, achieving the highest accuracy score as InternLM 2, excelling in PubMedQA Answer Verification and HealthFact Verification while in EMBS, MMedIns-Llama 3 slightly falls behind the GPT-4 with 11.99/12.90 in BLEU and ROUGE, which we treat as future work for further improvement.
[] Method PubMedQA PUBLICHEALTH EMBS MedNLI textual entailment Answer Ver. Health Fact Ver. Justification Ver. Discriminative Task Generative Task Close-source Models GPT-4 52 76 16.36/16.33 82 27.09/23.71 Claude-3.5 4 64 14.22/15.82 86 17.73/16.70 Open-source Models MEDITRON 26 40 10.86/13.73 66 3.92/14.27 InternLM 2 98 92 8.54/15.10 84 25.42/24.91 Llama 3 94 64 12.96/14.37 62 33.69/27.28 Mistral 54 68 15.19/15.56 72 20.81/19.03 MMedIns-Llama 3 98 92 11.99/12.90 84 37.56/32.1
Table 8: Results on fact verification and NLI results, as reported with both accuracy and BLUE/ROUGE scores. ‘Ver.’ denotes ‘verification’.
@@
InternLM 2 achieves the highest scores among the open-source LLMs, scoring 84. For the closed-source LLMs, GPT-4 and Claude-3.5 all show relatively high scores, with 82 and 86 accuracy scores respectively. In the generative task, Llama 3 shows the highest consistency with the referenced ground truth with 33.69/27.28 for BLEU and ROUGE. Mistral and Llama 3 exhibit mid-range performance, with Mistral achieving 72 accuracy and BLEU/ROUGE scores of 20.81/19.03. GPT-4 follows it, resulting in 27.09/23.71 scores while Claude-3.5 is not ideal in the generative task.
MMedIns-Llama 3 achieves the highest accuracy in the discriminative task, scoring 84, while slightly weaker than Claude-3.5. MMedIns-Llama 3 also excels in the generative task, with BLEU/ROUGE scores of 37.56/32.17, significantly outperforming other models in generating accurate and relevant entailments.
@@ 4.1 Data Collection
Figure 3: This pipeline of our method. a The data collection pipeline. We mainly collect data through filtering natural instructions and prompting well-organized BioNLP datasets. b The training and evaluation pipeline for our model leveraging the collected MedS-Ins. We leverage the instruction tuning training method to combine different datasets and evaluate the final model on multiple benchmarks comprehensively.
@@
Through the above procedure, we prompt an extra 85 tasks into a unified free-form question-answering format, combined with the filtered data, resulting in a totaling 13.5 million high-quality samples, covering 122 tasks, termed as MedS-Ins (the detailed 122 task information can be found in Supplementary Sec. A.3), which has shown to significantly improve the LLMs on clinical tasks.
@@ 4.2 Model Training
Few-shot Prompting. Here, the $I$ contains the few-shot examples, that allow the model to learn the input-output mapping on the fly. We simply obtain such instructions by randomly sampling other cases from the training set of the same task, and organizing them using a straightforward template as follows:: {mdframed}[backgroundcolor=gray!20] Case1: Input: {CASE1_INPUT}, Output: {CASE1_OUTPUT}
@@ 4.3 Baselines
  • MEDITRON [5] is a large-scale medical LLM with 70B parameters, further pre-trained on Llama 2. It leverages 21.1M medical papers, guidelines for further pre-training, and supervised finetuning on different MCQA datasets with context and chain-of-thought prompt styles.
@@ A.2 Task Category Details
In this paper, we conclude 19 medical-related task categories in total. For evaluation, we adopt 11 categories, covering 39 datasets, as shown by Supplementary Table 9. Notably, as the evaluation involves commercial models, for example, GPT-4 and Claude 3.5, it is extremely costly to adopt the original large-scale test split. Therefore, for some benchmarks (as will be detailed in the following), we randomly sampling a number of test cases. For the MCQA tasks, we use all of them, as the close-source models have been evaluated by other papers.
@@
Text Summarization. Text summarization denote a generative task, where a model is tasked to extract key information from extensive pieces of text. This task is particularly prevalent in the clinical field, for example, quickly extracting essential information from lengthy examination reports or helping clinicians and patients better understand complex medical reports. We evaluate the model’s performance using the following benchmarks.
@@
  • RCT-Text [44] is a dataset designed to summarize medical evidence from various clinical studies as presented in literature reviews. For this task, we provide the LLM with the titles and abstracts of studies, and task the model to output the primary conclusions of each study.
@@
  • MIMIC-IV Report [17] is a sub-set of MIMIC-IV. Specifically, MIMIC-IV is a large deidentified medical dataset of patients admitted to the emergency department or an intensive care unit at the Beth Israel Deaconess Medical Center in Boston, MA. It contains data for more than 65K patients admitted to an ICU and over 200K patients admitted to the emergency department. In this large-scale medical informatics database, there is a radiology note split that contains 2,321,355 deidentified radiology reports for 237,427 patients. For each radiology report, inspired by existing works [42, 50], we treat the impression part as a general summarization of the findings. We randomly sampled 100 cases for each covered modality and body region part, like Chest CT, Brain MRI, etc..
@@
  • PMC-patient basic information extraction [51] is a pioneering collection that includes 167,000 patient summaries, extracted from case reports in PubMed Central (PMC). This dataset has been further enriched with annotations detailing basic patient information. For this specific task, we provide the Large Language Model (LLM) with an extensive clinical text about the patient, and require the model to accurately extract information regarding the patient’s gender and age. During test, we randomly sample 100 instances for evaluation.
@@
  • PUBHEALTH Explanation [19] is a comprehensive dataset geared towards explainable automated fact-checking of public health claims. The task requires models to furnish an explanation for a specified claim using supporting material from the provided paragraph. During testing, we randomly sample 50 instances for evaluation.
  • HumanDiseaseOntology (DO) [36] is a database providing the biomedical community with consistent, reusable, and sustainable descriptions of human disease terms, phenotype characteristics, and related medical vocabularies. In this task, we ask the model to explain the specified given medical professional entity, and the description from the database functions as the ground truth. During testing, we randomly sample 100 instances for evaluation.
  • BioLORD Explanation [33] comprises pairs of biomedical concepts names and descriptions. In this task, we challenge models to elaborate on concise concepts by generating long, detailed definitions. During testing, we randomly sample 50 instances for evaluation.
  • MMedBench-Rationale [32] is a multilingual medical multiple-choice question and answering benchmark along with rationale for each sample. In this task, we will give the LLMs a question and its answer and require it to generate an explanation statement or step-by-step reasoning. During testing, we randomly sample 50 instances for each language.
@@
  • BC4CHEMD [35] is a dataset comprising 10,000 PubMed abstracts with a total of 84,355 chemical entity mentions, manually annotated by expert chemistry literature curators. In this task, the LLM is to input an abstract, and is required to recognize the name of the chemical. During testing, we randomly sample 50 instances for evaluation.
  • BC5CDR [22] is a widely-used resource in biomedical natural language processing, annotated for chemical and disease entities and their relationships. This benchmark is split into 2 task according to the category of the entity. In the first category, the LLM is required to recognize the name of the chemical. In the other category, the LLM is required to write the name of the disease. During testing, we randomly sample 50 instances for each task.
  • Species800 [29] comprises 800 PubMed abstracts in which organism mentions were identified. In this task, the LLM is input an abstract and is required to recognize the name of the organism. During testing, we randomly sample 50 instances for evaluation.
@@
  • DDXPlus [8] is a pioneering large-scale dataset tailored for Automatic Symptom Detection (ASD) and Automatic Diagnosis (AD) systems in the medical domain, featuring synthesized patient data from a proprietary medical knowledge base and commercial AD system. In this task, the model must make diagnostic decisions based on dialogues, selecting from a provided list of potential diagnoses. During testing, we randomly sample 50 instances for evaluation.
@@
  • SEER [7] is a treatment planning dataset leveraging the Surveillance, Epidemiology, and End Results (SEER) custom breast cancer databases. It collects 32,922 patient records and for each record with 19 attributes where the recommended treatment plans are noted, commonly in five types, i.e., Surgery Chemotherapy, Hormonal therapy, Biological therapy, and Radiation therapy. During testing, we randomly sample 50 instances for evaluation.
@@
  • MIMIC4ED Benchmark [49] provides a standardized benchmark derived from the Medical Information Mart for Intensive Care IV-Emergency Department (MIMIC-IV-ED) database, including data covering several key prediction tasks within the emergency department. We use this dataset for the comparison of LLMs for predicting clinical outcomes in emergency medicine. According to the differential of the categories of the clinical output, we split it into 3 tasks. We give the model an EHR of a patient and require the model to predict whether the patient needs, or the patient may revisit the emergency department in 72 hours, or the patient should be classified into a critical triage queue, separately. During testing, we randomly sample 100 instances for each task.
@@
  • HoC [1] is a specialized dataset containing 1,852 PubMed publication abstracts, which are expertly annotated according to a taxonomy. In this task, LLM is required for you to classify the hallmarks of cancer according to the given biomedical publication abstracts. During testing, we randomly sample 50 instances for evaluation.
@@
  • PUBHEALTH Verification [19] is a dataset designed for explainable automated fact-checking of public health claims. Beyond asking models to explain the fact details, it can also be modified as a fact verication task, i.e., given a paragraph along with a related claim and models are required to determine whether the claim contradicts the evidence provided in the paragraph. During testing, we randomly sample 25 instances for evaluation.
  • PubMedQA Verication [14] is a question-answering medical dataset based on PubMed abstracts. Unlike direct use in a Question Answering benchmark, we provide the LLM with a passage accompanied by a question and its answer. The Task for LLM is to verify if the answer answers the question. During testing, we randomly sample 50 instances for evaluation.
  • EBMS justification verification [27] During testing, we randomly sample 140 instances for evaluation.
Natural Language Inference (NLI). Natural Language Inference (NLI) is one of the critical tasks for understanding natural language. The objective of NLI is to determine if a given hypothesis can be inferred from a given premise. In medical this task is useful for making inferences form the textual medical history of patients.
  • MedNLI [34] is a nature language inference dataset for clinical domain, which is grounded in patient medical histories and annotated by medical professionals. We have divided the dataset into two tasks: ‘Entailment (discriminative)’ and ‘Entailment (generative)’. In the discriminative task, the model is initially presented with a formal clinical premise, such as condition descriptions or quantitative results. It is then given a hypothesis statement and must determine whether the hypothesis can be logically inferred from the clinical condition. In the generative task, the model is presented with a formal clinical premise and is required to generate a coherent and contextually correct hypothesis statement based on the given premise. During testing, we randomly sample 50 instances for each task.
@@ A.2.1 The Detailed Statistic of Our Benchmark
Multi-choice Question Answering MedQA, PubMedQA, MedMCQA, MMedBench Recalling facts from Pre-training 144 15 21 180 Accuracy
@@
Fact Verification PUBHEALTH Explanation, PubMedQA Verication Recalling facts from Pre-training 315 27 47 389 Accuracy
@@
Text Summarization PubMed Summarization, MedQSum, MIMIC-IV Report Retrieving facts from Contexts 372 79 20 470 BLEU, ROGUE
Information Extraction ADE Corpus, PICO Retrieving facts from Contexts 137 7 126 268 BLEU, ROGUE
Named Entity Recognition NCBI Disease Corpus, JNLPBA, BC2GM, BC4CHEMD, BC5CDR, Species80, ChemProt, Linnaeus, AnEM Retrieving facts from Contexts 35 5 52 92 Accuracy, Recall, Precision, F1

### v2

@@ 2.3 Quantitative Results on Various Tasks
InternLM 2 achieves the highest accuracy on PubMedQA Answer Verification with scores of 99.23. On PUBLICHEALTH, it also achieves 76.94 accuracy scores, which is the best among all the open-sources models, just behind GPT-4 with 78.60. In the EBMS benchmark, GPT-4 demonstrates the strongest performance, with BLEU/ROUGE scores of 16.28/16.27. Llama 3 achieves an accuracy of 94.77 and 63.89 on PubMedQA and HealthFact Verification, respectively, and a BLEU/ROUGE score of 16.52/16.49 in the EBMS benchmark. Mistral, while achieving a high BLEU/ROUGE score of 15.98/16.43 in EMBS, has a relatively lower accuracy of 57.38 and 69.78. MEDITRON ranks lowest across the accuracy metrics, with accuracy of 25.23 and 32.66 on the PubMedQA and PUBLICHEALTH datasets. In EMBS, MEDITRON achieves a BLEU/ROUGE score of 11.58/15.78.
MMedIns-Llama 3 continues surpassing existing models, achieving the highest accuracy score as InternLM 2, excelling in PubMedQA Answer Verification and HealthFact Verification while in EMBS, MMedIns-Llama 3 slightly falls behind the GPT-4 and Llama 3 with 12.71/14.65 in BLEU and ROUGE, which we treat as future work for further improvement.
[] Method PubMedQA PUBLICHEALTH EMBS MedNLI textual entailment Answer Ver. Health Fact Ver. Justification Ver. Discriminative Task Generative Task Close-source Models GPT-4 66.15 78.60 16.28/16.27 86.63 27.09/23.71 Claude-3.5 11.54 62.04 14.77/16.45 82.14 17.80/20.02 Open-source Models MEDITRON 25.23 32.66 11.58/15.78 60.83 4.42/14.08 InternLM 2 99.23 76.94 8.75/14.69 84.67 15.84/19.01 Llama 3 94.77 63.89 16.52/16.49 63.85 21.31/22.75 Mistral 57.38 69.78 15.98/16.43 71.59 13.03/15.47 MMedIns-Llama 3 97.08 79.55 12.71/14.65 86.71 23.52/25.17
Table 8: Results on fact verification and NLI results, as reported with both accuracy and BLUE/ROUGE scores. ‘Ver.’ denotes ‘verification’. Bolding represents the best results.
@@
InternLM 2 achieves the highest scores among the open-source LLMs, scoring 84.67. For the closed-source LLMs, GPT-4 and Claude-3.5 all show relatively high scores, with 86.63 and 82.14 accuracy scores respectively. Mistral shows the second best accuracy score of 71.59 among the open-source LLMs. In the generative task, Llama 3 shows the highest consistency with the referenced ground truth with 21.31/22.75 for BLEU and ROUGE among the open-source models. Mistral and InternLM 2 exhibit mid-range performance in the generative task, with Mistral achieving BLEU/ROUGE scores of 13.03/15.47, InternLM 2 achieving 15.84/19.01 scores. Similarly, GPT-4 also performs well in the generative task format, resulting in 27.09/23.71 scores while Claude-3.5 is not ideal in this task.
MMedIns-Llama 3 achieves the highest accuracy in the discriminative task, scoring 86.71, comparable with GPT-4. MMedIns-Llama 3 also excels in the generative task, with BLEU/ROUGE scores of 23.52/25.17, outperforming other models except the GPT-4.
@@ 4.1 Data Collection
Figure 3: The pipeline of our method. a The data collection pipeline. We mainly collect data through filtering natural instructions and prompting well-organized BioNLP datasets. b The training and evaluation pipeline for our model leveraging the collected MedS-Ins. We leverage the instruction tuning training method to combine different datasets and evaluate the final model on multiple benchmarks comprehensively.
@@
Through the above procedure, we prompt an extra 85 tasks into a unified free-form question-answering format, combined with the filtered data, resulting in a totaling 5M instances with 19K instructions, covering 122 tasks, termed as MedS-Ins (the detailed 122 task information can be found in Supplementary Sec. A.3), which has shown to significantly improve the LLMs on clinical tasks.
@@ 4.2 Model Training
Few-shot Prompting. Here, the $I$ contains the few-shot examples, that allow the model to learn the input-output mapping on the fly. We simply obtain such instructions by randomly sampling other cases from the training set of the same task, and organizing them using a straightforward template as follows: {mdframed}[backgroundcolor=gray!20] Case1: Input: {CASE1_INPUT}, Output: {CASE1_OUTPUT}
@@ 4.3 Baselines
  • MEDITRON [5] is a large-scale medical LLM with 7B parameters, further pre-trained on Llama 2. It leverages 21.1M medical papers, guidelines for further pre-training, and supervised finetuning on different MCQA datasets with context and chain-of-thought prompt styles.
@@ A.2 Task Category Details
In this paper, we conclude 19 medical-related task categories in total. For evaluation, we adopt 11 categories, covering 28 datasets, as shown by Supplementary Table 9. To ensure the integrity of our evaluation, we deliberately exclude the test set from the source dataset during the prompting of instances. This test set is subsequently utilized as a benchmark for assessing the performance of LLMs. For datasets that do not have a predefined train-test split, we randomly partition the dataset into training and testing subsets at an 9:1 ratio. Notably, as the evaluation involves commercial models, for example, GPT-4 and Claude 3.5, it is extremely costly to adopt the original large-scale test split. Therefore, for some benchmarks (as will be detailed in the following), we randomly sampling a number of test cases. For the MCQA tasks, we use all of them, as the close-source models have been evaluated by other papers.
@@
Text Summarization. Text summarization denotes a generative task, where a model is tasked to extract key information from extensive pieces of text. This task is particularly prevalent in the clinical field, for example, quickly extracting essential information from lengthy examination reports or helping clinicians and patients better understand complex medical reports. We evaluate the model’s performance using the following benchmarks. Notably, different from other tasks, text summarization commonly involves longer contexts which is more costly for evaluating the close-source models, thus we adopt smaller sampling cases (100) for evaluation.
@@
  • RCT-Text [44] is a dataset designed to summarize medical evidence from various clinical studies as presented in literature reviews. For this task, we provide the LLM with the titles and abstracts of studies and task the model to output the primary conclusions of each study. During testing, we randomly sample 100 instances for evaluation.
@@
  • MIMIC-IV Report [17] is a sub-set of MIMIC-IV. Specifically, MIMIC-IV is a large deidentified medical dataset of patients admitted to the emergency department or an intensive care unit at the Beth Israel Deaconess Medical Center in Boston, MA. It contains data for more than 65K patients admitted to an ICU and over 200K patients admitted to the emergency department. In this large-scale medical informatics database, there is a radiology note split that contains 2,321,355 deidentified radiology reports for 237,427 patients. For each radiology report, inspired by existing works [42, 50], we treat the impression part as a general summarization of the findings. We randomly sampled 100 cases for each covered modality and body region part, like Chest CT, Brain MRI, etc., resulting in 1430 cases (for some parts, the test split has fewer than 100 cases, we directly keep all.) for evaluation, covering different anatomies and modalities.
@@
  • PMC-patient basic information extraction [51] is a pioneering collection that includes 167,000 patient summaries, extracted from case reports in PubMed Central (PMC). This dataset has been further enriched with annotations detailing basic patient information. For this specific task, we provide the Large Language Model (LLM) with an extensive clinical text about the patient, and require the model to accurately extract information regarding the patient’s gender and age. During test, we randomly sample 1500 instances for evaluation.
@@
  • PUBHEALTH Explanation [19] is a comprehensive dataset geared towards explainable automated fact-checking of public health claims. The task requires models to furnish an explanation for a specified claim using supporting material from the provided paragraph. During testing, we use the entire test set consisting of 1235 samples.
  • HumanDiseaseOntology (DO) [36] is a database providing the biomedical community with consistent, reusable, and sustainable descriptions of human disease terms, phenotype characteristics, and related medical vocabularies. In this task, we ask the model to explain the specified given medical professional entity, and the description from the database functions as the ground truth. During testing, we use the entire test set of the dataset, which consists of 1115 instances.
  • BioLORD Explanation [33] comprises pairs of biomedical concepts names and descriptions. In this task, we challenge models to elaborate on concise concepts by generating long, detailed definitions. During testing, we use the entire test set of the dataset, which consists of 1115 instances.
  • MMedBench-Rationale [32] is a multilingual medical multiple-choice question and answering benchmark along with rationale for each sample. In this task, we will give the LLMs a question and its answer and require them to generate an explanation statement or step-by-step reasoning. During testing, We utilize the entire test set of the dataset, comprising detailed allocations of samples for each language: 200 for Chinese, 195 for English, 189 for French, 190 for Japanese, 173 for Russian, and 189 for Spanish.
@@
  • BC4Chem [35] is a dataset comprising 10,000 PubMed abstracts with a total of 84,355 chemical entity mentions, manually annotated by expert chemistry literature curators. In this task, the LLM is to input an abstract, and is required to recognize the name of the chemical. During testing, we randomly sample 1500 instances for evaluation.
  • BC5CDR (BC5Chem and BC5Disease) [22] is a widely-used resource in biomedical natural language processing, annotated for chemical and disease entities and their relationships. This benchmark is split into 2 task according to the category of the entity. In the first category, the LLM is required to recognize the name of the chemical. In the other category, the LLM is required to write the name of the disease. During testing, we randomly sample 1500 instances for each task.
  • Species800 [29] comprises 800 PubMed abstracts in which organism mentions were identified. In this task, the LLM is input an abstract and is required to recognize the name of the organism. During testing, we randomly sample 1500 instances for evaluation.
@@
  • DDXPlus [8] is a pioneering large-scale dataset tailored for Automatic Symptom Detection (ASD) and Automatic Diagnosis (AD) systems in the medical domain, featuring synthesized patient data from a proprietary medical knowledge base and commercial AD system. In this task, the model must make diagnostic decisions based on dialogues, selecting from a provided list of potential diagnoses. During testing, we randomly sample 1500 instances for evaluation.
@@
  • SEER [7] is a treatment planning dataset leveraging the Surveillance, Epidemiology, and End Results (SEER) custom breast cancer databases. It collects 32,922 patient records and for each record with 19 attributes where the recommended treatment plans are noted, commonly in five types, i.e., Surgery Chemotherapy, Hormonal therapy, Biological therapy, and Radiation therapy. During testing, we randomly sample 1500 instances for evaluation.
@@
  • MIMIC4ED Benchmark [49] provides a standardized benchmark derived from the Medical Information Mart for Intensive Care IV-Emergency Department (MIMIC-IV-ED) database, including data covering several key prediction tasks within the emergency department. We use this dataset for the comparison of LLMs for predicting clinical outcomes in emergency medicine. According to the differential of the categories of the clinical output, we split it into 3 tasks. We give the model an EHR of a patient and require the model to predict whether the patient needs, or the patient may revisit the emergency department in 72 hours, or the patient should be classified into a critical triage queue, separately. During testing, we randomly sample 1500 instances for each task.
@@
  • HoC [1] is a specialized dataset containing 1,852 PubMed publication abstracts, which are expertly annotated according to a taxonomy. In this task, LLM is required for you to classify the hallmarks of cancer according to the given biomedical publication abstracts. During testing, we use the entire test set of the dataset, which consists of 158 instances for each task.
@@
  • PUBHEALTH Verification [19] is a dataset designed for explainable automated fact-checking of public health claims. Beyond asking models to explain the fact details, it can also be modified as a fact verication task, i.e., given a paragraph along with a related claim and models are required to determine whether the claim contradicts the evidence provided in the paragraph. During testing, we use the entire test set of the dataset, which consists of 650 instances for each task.
  • PubMedQA Verication [14] is a question-answering medical dataset based on PubMed abstracts. Unlike direct use in a Question Answering benchmark, we provide the LLM with a passage accompanied by a question and its answer. The Task for LLM is to verify if the answer answers the question. During testing, we use the entire test set of the dataset, which consists of 1188 instances for each task.
  • EBMS justification verification [27] During testing, we use the entire test set of the dataset, which consists of 304 instances for each task.
Natural Language Inference (NLI). Natural Language Inference (NLI) is one of the critical tasks for understanding natural language. The objective of NLI is to determine if a given hypothesis can be inferred from a given premise. In medical this task is useful for making inferences from the textual medical history of patients.
  • MedNLI [34] is a nature language inference dataset for clinical domain, which is grounded in patient medical histories and annotated by medical professionals. We have divided the dataset into two tasks: ‘Entailment (discriminative)’ and ‘Entailment (generative)’. In the discriminative task, the model is initially presented with a formal clinical premise, such as condition descriptions or quantitative results. It is then given a hypothesis statement and must determine whether the hypothesis can be logically inferred from the clinical condition. In the generative task, the model is presented with a formal clinical premise and is required to generate a coherent and contextually correct hypothesis statement based on the given premise. During testing, we use the entire test set of the dataset, which consists of 1422 instances for each task.
@@ A.2.1 The Detailed Statistic of Our Benchmark
Multi-choice Question Answering MedQA, PubMedQA, MedMCQA, IgakuQA, RuMedBench, FrenchMedMCQA, HeadQA Recalling facts from Pre-training 144 15 21 180 Accuracy
@@
Fact Verification PUBHEALTH Explanation, PubMedQA Verication, EBMS Recalling facts from Pre-training 315 27 47 389 Accuracy
@@
Text Summarization MIMIC-IV, RCT, MIMIC-CXR, MedQSum Retrieving facts from Contexts 372 79 20 470 BLEU, ROGUE
Information Extraction ADE, PICO, PMC-Patient Retrieving facts from Contexts 137 7 126 268 BLEU, ROGUE
Named Entity Recognition BC4Chem, BC5Chem, BC5Disease, Species800 Retrieving facts from Contexts 35 5 52 92 Accuracy, Recall, Precision, F1